# FrameFusion: Combining Similarity and Importance for Video Token Reduction on Large Visual Language Models

### v1

@@ 4 FrameFusion Design
Figure 6: FrameFusion first merges tokens with similarities above a specified threshold at shallow layers, then applies top-$k$ importance pruning to comply with the given computational constraints.
Building upon the observations from Section 3, we introduce FrameFusion, a novel token compression technique for video LVLMs, exploring a new perspective of token similarity. The design of FrameFusion is detailed in Section 4.1, followed by an explanation of the rationale behind each key design choice in Section 4.2.
@@ 4.1 Two-Stage Token Compression
The core concept of FrameFusion is illustrated in Figure 1: unlike traditional methods that primarily employ importance-based token pruning, FrameFusion also emphasizes similarity-based token merging. It merges similar tokens and prunes unimportant ones, retaining only those that are both important and unique. The two-stage token reduction process of FrameFusion is shown in Figure 6.
Merging stage. In the initial merging stage, FrameFusion utilizes token similarity to merge vision tokens. It computes the token similarity $S^{(l)}$ at the current layer using Equation 1, where only $N$ cosine similarities are computed between the corresponding vision tokens of adjacent frames. Tokens that exceed the similarity threshold $S_{\text{threshold}}$ are grouped with their analogous tokens from the previous frame. These merging groups are transitive, allowing concatenated groups to form a larger group containing more than two tokens as shown in Figure 6. Within each group, FrameFusion performs element-wise averaging of all tokens, assigning the result to the first token in the group. This forward merging strategy ensures that subsequent vision tokens can still aggregate information from all preceding tokens using causal attention.
The merging stage is applied across successive shallow LLM layers to progressively merge similar tokens, until the number of similar tokens falls below the threshold $N_{\text{threshold}}$ at a specific layer. Merging occurs right before the FFN module; however, to reduce the number of tokens for the attention module at the first LLM layer, an additional merging step is performed before it. After the merging stage, the remaining unique tokens advance to the pruning stage.
Pruning stage. After the merging stage, FrameFusion further prunes unimportant tokens. As discussed in Section 3.1, FrameFusion uses cumulative attention scores to represent token importance. Based on a user-defined computational cost budget, FrameFusion calculates the maximum number of remaining tokens $k$ that fits within the budget. It then applies top-$k$ importance pruning to retain only the important tokens from the remaining unique ones.
Through the merging and pruning stages, only the unique and important vision tokens are retained for subsequent processing, significantly boosting LVLM efficiency.
@@ 4.2 Design Choice Rationales
In this subsection, we explain the key design choices of FrameFusion with rationales from key observations in Section 3.
Unlike token importance, which reuses the existing $N\times N$ attention scores, token similarity introduces a new, orthogonal metric. To avoid additional $N\times N$ similarity computations for all token pairs, we leverage Observation 2 to compute only empirically similar token pairs with an $O(N)$ complexity:
Design Choice 1. FrameFusion computes token similarities only between corresponding vision tokens of adjacent frames.
Another key design choice is determining the layers at which to perform token reduction. For importance-based pruning, previous study [2] observes the reduction in vision token importance after the initial two layers for LVLMs. Thus, it advises applying importance-based pruning after the second layer. Other works also confirm the accuracy advantage of aggressive pruning only at deeper layers [8, 1]. In contrast, merging relies on high token similarities and therefore prefers shallow layers, according to Observation 3. Based on the contradictory preferences of similarity-based merging and importance-based pruning, FrameFusion is designed as follows:
Design Choice 2. FrameFusion applies token merging at the initial successive layers, then pruning at a later layer.
The final key design choice is whether merged tokens should remain reduced in subsequent layers, i.e., whether to use cascaded token merging. This choice depends on whether tokens merged in shallow layers are likely to be needed in deeper layers. Given the high rank consistency and retention rate of token similarities discussed in Observation 4, we adopt the following design to achieve a better efficiency trade-off:
@@
These rationales for the three key design choices in FrameFusion enhance the explainability of its better performance.
@@ 5.1 Setups
Baselines. We compare FrameFusion with state-of-the-art token pruning baselines, StreamingLLM [26] and FastV [2]. Hyperparameters follow the official implementations and are detailed in Appendix7.
Token Budget. We term the token budget as the average number of tokens required for the KV-Cache at the start of the decoding stage. The relative token budget, denoted as $C$, is the token budget of compression methods divided by the original input length $N$. The relative token budget approximates the relative KV-Cache utilization for each compression method. We set $C=30\%$ for all token compression methods if not specified. Specifically, for the cascaded methods FrameFusion and FastV, the token budget equals the average token length per layer across both the prefill and decode stages. For StreamingLLM, the token budget equals the token length in KV-Cache at the decode stage, which also equals the attention span during the prefill stage.
Models. We evaluate our approach on widely used video LVLMs, including lmms-lab models: LLaVA-Video-{7B,72B}-Qwen2 (referred to as Llava-Video-{7B,72B}), LLaVA-NeXT-Video-32B-Qwen (referred to as Llava-Video-32B) [34], and the openbmb model MiniCPM-V-2_6 (referred to as MiniCPM-V-8B) [29].
Benchmarks. We use lmms-eval [32] as the primary evaluation framework to test diverse video benchmarks, including VideoMME [6] for video understanding, NExT-QA [27] for video question-answering, and Video Needle In A Haystack (VideoNIAH)[36] for visual content retrieval.
5.2 Accuracy-Computation Trade-off
@@
As shown in Figure 7, we explore the accuracy-computation trade-off of different token compression methods. The relative FLOPs is normalized by the original dense model with the video sampling rate at 1 frame per second. For the original variant, the relative FLOPs is altered by directly multiplying the sampling rate with the relative token budget. FrameFusion advances the Pareto Front in video content retrieval accuracy and computing FLOPs across varied relative token budgets.
@@ 5.3 Performance
VideoMME $\uparrow$ NExt-QA $\uparrow$ VideoNIAH $\uparrow$
Model Method Short Medium Long Multi Choice Open End Edit Insert1 Insert2 Avg.
Llava-Video-7B Original 75.8 61.7 52.2 83.2 32.1 90.7 50.7 88.0 66.8
StreamingLLM 63.4 54.1 46.4 79.0 30.3 26.0 15.3 28.7 42.9
FastV 68.4 58.0 49.6 81.1 31.2 69.3 28.7 76.7 57.9
Ours 74.0 59.8 50.0 81.8 31.7 90.0 48.7 87.3 65.4
MiniCPM-V-8B Original 69.1 56.6 49.8 78.9 13.8 88.7 36.7 88.7 60.3
StreamingLLM 61.1 51.8 48.4 76.0 23.2 22.0 15.3 28.7 40.8
FastV 67.1 53.9 49.2 78.0 14.8 82.7 26.7 71.3 55.5
Ours 69.7 54.1 48.3 78.2 16.3 89.3 41.3 89.3 60.8
Table 1: The overall performance comparison across different models and methods on VideoMME, NextQA, and VideoNIAH benchmarks. All token compression methods employ 30% relative token budget.
FrameFusion outperforms state-of-the-art token compression methods across various models and benchmarks, achieving performance comparable to the dense model at a 30% token budget (Table 1). Additional results for additional budgets are in Appendix 8.1. Across VideoMME, NExT-QA, and VideoNIAH benchmarks, FrameFusion limits relative performance drops to under 3%. In contrast, StreamingLLM and FastV show drops of up to 13% and 7% for VideoMME and NExT-QA, and 70% and 24% for VideoNIAH. Further retrieval performance details are in Appendix 8.2.
@@ 5.4 Efficiency
We evaluate efficiency on a single NVIDIA A100-80GB GPU, averaging results over 128 videos from the VideoMME benchmark. Figure 8 highlights how FrameFusion improves runtime and memory efficiency for the colored components.
Time. FrameFusion achieves 3.4-4.4$\times$ speedups with a 30% token budget across various input frames. Including other runtimes like ViT and CPU calls, FrameFusion maintains 1.6 – 1.9$\times$ end-to-end speedups. Speedup results for additional model sizes are provided in Appendix 8.3.
Memory. FrameFusion reduces LLM memory consumption for KV-Cache and corresponding activations (noted as “KV-Cache” in Figure 8) to 44% – 51% with a 30% token budget, enabling more frames to fit within GPU memory limits.
5.5 Scalability
Scaling Model Size.
We test the scalability of FrameFusion across diverse model parameter sizes with the Llava-Video model family. As shown in Figure 13, FrameFusion consistently outperforms FastV baseline across all model sizes and VideoMME categories, demonstrating comparative performance with the original model at a 30% relative token budget. The numeric results are shown in Appendix 8.5.
Figure 10: The VideoNIAH performance for the Llava-Video-7B across various numbers of input frames.
Scaling Input Length. Figure 10 shows that FrameFusion surpass other baselines to achieve high retrieval accuracies across diverse input number of frames, demonstrating comparable performance with the original dense model. Figure 8 shows increasing LLM speedups and memory reductions with more input frames.
@@ 5.6 Ablation Study
We validate the impact of each observation and design choice in FrameFusion, demonstrating their individual effectiveness in this section.
Figure 11: Top-k hit numbers for different similarity calculation strategies on the initial layers of the Llava-Video-7B model.
Design Choice 4.2. Instead of computing $N\times N$ token similarities with significant overhead, FrameFusion calculates $N$ token similarities only between corresponding vision tokens in adjacent frames. We compare our approach with two common alternatives: 1. The adjacent strategy, which computes $N$ similarities between adjacent image patches (i.e., adjacent vision tokens). 2. The random strategy, which calculates similarities for $N$ randomly selected token pairs. All three strategies are compared against the posterior optimal upper bound, which merges the most similar tokens using the full $N\times N$ similarity computation. As shown in Figure 11, we evaluate the top-K hit number for each strategy. This metric indicates how many of the top-k similar token pairs identified by a given strategy align with those selected by the posterior optimal approach. For visualization clarity, we limit the number of frames to three (420 non-first-frame vision tokens) from each video. Our method consistently outperforms the adjacent and random strategies across all model layers.
Merging Layer Budget Merging Rate Retrieval Acc.
Original 100% 0% 76.4%
0 50% 50.0% 76.2%
1 50% 52.0% 76.8%
2 50% 53.8% 76.4%
12 50% 87.5% 74.4%
13 50% 93.3% 64.2%
14 50% 100.0% 48.9%
@@
Table 2: Performance on VideoNIAH of different merging layers with the same relative token budget of 50%.
Method VideoMME $\uparrow$ VideoNIAH $\uparrow$
Original 63.2 76.4
Prune then merge 59.9 73.1
Merge then prune 60.9 73.3
@@
Table 3: Performance of the Llava-Video-7B model of different orders between merging and pruning.
@@
Design Choice 4.2. FrameFusion first merges at the initial layers, then prunes at subsequent layers. We validate the influence of the merging layer position. As shown in Table 2, given the same token reduction budget, merging at shallow layers shows significantly better performance than merging at deeper layers. Additionally, we examine the effect of the order of merging and pruning. As shown in Table 3, with a fixed token budget of $C=0.3$ and the same number of tokens reduced in layers 1 and 2, merging before pruning achieves better performance compared to pruning before merging.
[…]
#Frame w. Cascaded Original Ours (Cascaded)
8 92.8% 20.0 20.7
16 90.0% 29.1 28.0
32 81.5% 43.8 42.9
64 72.2% 76.4 75.6
Table 4: Equivalent relative token budgets and VideoNIAH scores of cascaded merging across different numbers of frames, aligned to the minimum FLOPs achievable by non-cascaded merging.
@@
Design Choice 4.2. FrameFusion adopts a cascaded merging strategy, where merged tokens remain reduced across layers to maximize efficiency. To evaluate the accuracy-efficiency trade-offs of cascaded merging, we use non-cascaded merging (KV-cache merging) as a baseline. This baseline reduces only the computation of attention matrices while leaving FFN computations unchanged. We calculate the minimum FLOPs achievable with non-cascaded merging by removing all KV-caches. As shown in Table 4, under the same FLOPs, FrameFusion requires token budgets of only 72% to 93% when using cascaded pruning—substantially more relaxed than the 30% budgets used in the main experiments. Across different numbers of frames, FrameFusion achieves performance comparable to the original model, while the non-cascaded model will fail due to the absence of KV-caches.

### v2

@@ 4 FrameFusion Design
Figure 6: FrameFusion first (a) merges tokens with similarities above a specified threshold at shallow layers, then (b) applies top-$k$ importance pruning to comply with the given computational constraints. (c) Tokens are permanently reduced for subsequent layers.
Building upon the observations from Section 3, we propose FrameFusion, a novel token compression method for video LVLMs, exploring the new perspective of token similarity. The detailed design is introduced in Section 4.1, followed by the rationales behind key design choices in Section 4.2.
@@ 4.1 Two-Stage Token Compression
The core idea of FrameFusion is illustrated in Figure 1. Unlike traditional methods that solely rely on importance-based pruning, FrameFusion additionally merges similar tokens before pruning, retaining only those that are both important and unique. The two-stage token compression approach is depicted in Figure 6.
Merging stage. In the first merging stage, FrameFusion utilizes token similarity to merge visual tokens. Specifically, it computes token similarity $S^{(l)}$ at each shallow layer according to Equation 1, considering only $N$ cosine similarities between corresponding tokens from adjacent frames. Tokens whose similarity exceeds a predefined threshold ($S_{\text{threshold}}$) are grouped with their corresponding tokens from previous frames. These merging groups are transitive, allowing concatenated groups to form a larger group containing more than two tokens (as shown in Figure 6). Within each group, FrameFusion performs element-wise averaging of all tokens and assigns the averaged result to the earliest token in the group. This forward merging strategy ensures that subsequent visual tokens can still aggregate information from all preceding tokens using causal attention.
This merging procedure is progressively applied at successive shallow LLM layers until the number of similar tokens falls below a predefined threshold ($N_{\text{threshold}}$). Merging is applied before the feed-forward network (FFN) module. However, to further reduce token numbers in the attention module of the first LLM layer, an additional merging step is performed before it. After completing the merging stage, the remaining tokens proceed to the pruning stage.
Pruning stage. After the merging stage, FrameFusion further prunes unimportant tokens. As defined in Section 3.1, FrameFusion uses cumulative attention scores to represent token importance. Given a user-specified computational budget, FrameFusion determines the maximum allowable number of tokens ($k$). It then applies pruning to retain only the top-$k$ tokens based on importance scores.
Through the merging and pruning stages, FrameFusion effectively retains only unique and important visual tokens for subsequent processing, significantly enhancing the efficiency of LVLMs. We also investigate the effect of different $S_{\text{threshold}}$ and $N_{\text{threshold}}$. Experimental results indicate that adjustments to these thresholds result in only minor variations in model performance. Detailed results are presented in Section 8.3.3.
@@ 4.2 Design Choice Rationales
In this subsection, we explain the rationales behind the key design choices of FrameFusion, grounded in the observations from Section 3. More empirical evidences are presented in Section 5.5.
Design Choice 1. FrameFusion computes token similarities only between corresponding visual tokens of adjacent frames.
Unlike token importance, which reuses the existing $N\times N$ attention scores, token similarity introduces a new, orthogonal metric. To avoid additional $N\times N$ similarity computations for all token pairs, we exploit Observation 2 to compute only empirically similar token pairs with an $O(N)$ complexity.
Design Choice 2. FrameFusion applies token merging at the initial successive layers, followed by pruning at deeper layers.
Another critical design choice involves determining the appropriate layers for different token reduction methods. For importance-based pruning, previous studies indicate a decline in visual token importance after the initial layers [3], recommending less pruning at shallow layers [9, 1]. In contrast, similarity-based merging depends on initially high token similarities, preferring shallow layers as per Observation 3. Given these contrasting preferences, FrameFusion employs merging at shallow layers and pruning at deeper layers to optimize both similarity and importance metrics.
@@
The final design choice addresses whether merged tokens should remain combined across subsequent layers (cascaded merging) or be individually reconsidered at each layer (non-cascaded merging).
Specifically, as shown in Figure 6(c), cascaded merging permanently reduces the token count once tokens are merged, significantly lowering computational costs in both feed-forward network (FFN) and attention modules at subsequent layers. In contrast, non-cascaded merging maintains the original token count at every layer. It selectively reduces computations within certain modules, typically pruning only the Key and Value matrices in attention layers [26, 20, 43]. Although non-cascaded merging retains flexibility by potentially reusing tokens at deeper layers, it incurs additional computational overhead due to repeated similarity evaluations and unchanged FFN computations.
Given these accuracy-efficiency trade-offs, the optimal merging strategy depends on whether tokens merged at shallow layers remain similar in deeper layers. Considering the higher consistency of token similarity rankings across layers (Observation 5), FrameFusion employs cascaded merging to eliminate unnecessary computations.
These rationales illustrate the motivation behind each design choice, clarifying the superior performance of FrameFusion.
@@ 5.1 Setups
Baselines. We compare FrameFusion with state-of-the-art token pruning baselines, StreamingLLM [31], FastV [3], and PruMerge [25]. Hyperparameters adhere to respective official implementations and are detailed in Appendix 7.
Models. We evaluate our approach across six video LVLMs from diverse model families and sizes, including lmms-lab models: LLaVA-Video-{7B,72B}-Qwen2 (denoted Llava-Video-{7B,72B}) [42]; NVLabs models: NVILA-Lite-2B, NVILA-8B-Video, NVILA-Lite-15B-Video (denoted NVILA-{2B,8B,15B}) [23]; and OpenBMB model MiniCPM-V-2_6 (denoted MiniCPM-V-8B) [34]. The PruMerge baseline is incompatible with MiniCPM-V due to its Q-Former architecture and is excluded for this model.
Benchmarks. We use lmms-eval [40] as the primary evaluation framework and test five video benchmarks: Video Needle In A Haystack (VideoNIAH)[44] for visual content retrieval; NExT-QA [32] for video question-answering; and VideoMME [7], EgoSchema [24] and MVBench [19] for general video understanding, highlighting spacial and temporal understanding, respectively.
Token Budget. We define token budget as the average sequence length of KV-Cache at the start of the decoding stage. For cascaded methods (FrameFusion, FastV, and PruMerge), it also equals the average token length per layer of the prefill stage. For StreamingLLM, it equals the sink size plus window size. The relative token budget, denoted $C$, is the token budget divided by the original input length $N$. Unless specified otherwise, we set $C=30\%$ for all token compression methods.
5.2 Computation-Accuracy Trade-off
@@
Figure 7 explores the computation-accuracy trade-offs of FrameFusion by varing the relative token budget. The x-axis shows the relative computation FLOPs, normalized to the original dense model operating at a 1 frame-per-second (fps) sampling rate. The y-axis shows the VideoNIAH retrieval accuracy. Higher accuracy at lower FLOPs (towards the top-left) indicates better trade-offs. The Original baseline, which directly adjusts the sampling rate of the original model, shows that FrameFusion achieves faster accuracy gains per FLOP compared to directly increasing frame rates. Other baselines also show significant accuracy degradation at reduced token budgets. In contrast, FrameFusion maintains high accuracy even at 30% computing FLOPs, greatly advancing the Pareto Front. We also explore the computation-accuracy trade-offs on different models and benchmarks. The results are detailed in Appendix 8.1.1.
@@ 5.3 Performance
Model Size Method VideoNIAH NExT-QA VideoMME EgoSchema MVBench Avg.$\uparrow$
edit insert1 insert2 mc oe w/o sub. w/ sub.
Llava-Video 7B Original 90.7 50.7 88.0 83.2 32.1 63.2 69.8 53.4 61.9 65.9
StreamingLLM 26.0 15.3 28.7 79.0 30.3 54.7 65.5 46.6 55.2 44.6
FastV 69.3 28.7 76.7 81.1 31.2 58.7 67.0 50.1 58.0 57.9
PruMerge 83.3 36.0 83.3 79.4 30.8 60.0 68.6 50.7 56.0 60.9
Ours 90.0 48.7 87.3 81.8 31.7 61.3 69.9 53.0 59.7 64.8
72B Original 89.3 66.0 88.0 85.3 32.3 70.9 77.3 65.0 63.9 70.9
StreamingLLM 33.3 20.0 35.3 81.9 30.6 62.6 72.9 60.2 58.0 50.5
FastV 22.0 48.7 77.3 83.7 31.5 65.9 73.7 62.6 61.7 58.6
PruMerge 85.3 58.0 86.0 82.0 31.4 66.7 74.8 62.6 58.6 67.3
Ours 90.0 63.3 88.0 84.6 32.0 69.0 76.7 63.2 63.0 70.0
NVILA 2B Original 90.0 22.0 87.3 71.2 6.6 50.9 53.2 42.3 50.7 52.7
StreamingLLM 26.0 12.7 34.7 69.0 5.8 45.7 50.1 40.7 49.1 37.1
FastV 50.7 14.7 56.7 70.7 7.2 46.7 50.6 41.1 50.1 43.2
PruMerge 27.3 31.3 81.3 67.7 11.1 47.3 50.4 42.2 48.0 45.2
Ours 89.3 27.3 87.3 71.8 20.1 50.4 53.1 45.2 49.5 54.9
8B Original 98.7 40.7 100.0 81.7 33.0 63.9 68.3 52.0 67.5 67.3
StreamingLLM 30.0 17.3 41.3 78.4 30.8 54.3 63.7 46.2 58.1 46.7
FastV 87.3 33.3 90.7 80.4 32.5 59.5 66.8 50.5 64.5 62.8
PruMerge 4.7 32.0 93.3 77.1 31.4 56.9 65.1 49.4 57.9 52.0
Ours 96.0 38.0 98.7 80.7 32.5 61.1 68.2 52.5 65.0 65.9
15B Original 95.3 42.0 100.0 78.7 30.9 65.8 72.3 58.2 60.5 67.1
StreamingLLM 34.0 18.7 34.0 74.0 28.5 58.5 65.1 53.7 55.0 46.8
FastV 48.7 24.7 80.7 77.0 30.6 60.6 69.1 56.7 57.3 56.2
PruMerge 19.3 43.3 98.0 72.4 30.0 59.3 68.4 52.3 52.8 55.1
Ours 94.0 52.7 99.3 77.7 31.2 63.5 70.8 57.8 58.4 67.3
MiniCPM-V 8B Original 88.7 36.7 88.7 78.9 13.8 58.5 60.3 53.4 55.0 59.3
StreamingLLM 22.0 15.3 28.7 76.0 23.2 53.8 56.7 48.2 51.3 41.7
FastV 82.7 26.7 71.3 78.0 14.8 56.7 58.2 51.8 53.2 54.8
Ours 89.3 41.3 89.3 78.2 16.3 57.4 59.5 52.3 53.6 59.7
Table 1: Performance comparison across different model families, sizes, and methods on five benchmarks at a 30% relative token budget.
Overall Performance. FrameFusion consistently outperforms state-of-the-art token compression methods across multiple model families, sizes, and benchmarks, matching dense model performance at a 30% token budget (Table 1). FrameFusion maintains a maximum relative average performance drop of just 2.4% across six models, whereas StreamingLLM, FastV, and PruMerge incur drops of 35.8%, 19.7%, and 15.4%, respectively. On VideoNIAH benchmark, FrameFusion shows a maximum relative drop of only 2.8%, compared to 38.7-69.5% for other methods. VideoMME is sensitive because it relies on spatial and temporal details, which cannot be answered by LLM common sense alone. Excluding VideoNIAH, FrameFusion’s maximum drop is 3.6%, significantly lower than the 8.3-14.2% seen in baselines. Detailed statistics are in Appendix 8.1.
Figure 8: The VideoNIAH performances for the Llava-Video-7B across various numbers of input frames.
Scaling Input Length. Figure 8 investigates model performance across varying frame counts, from 8 to 256 frames. FrameFusion consistently outperforms baseline methods, achieving retrieval accuracies close to the original dense model. As the frame count increases, the performance gap between FrameFusion and the dense model shrinks from 4.6% to 0.5%, significantly outperforming the best baseline, which has gaps ranging from 12.8% down to 5.8%. We also explore the scalability of FrameFusion across different token budgets, models and benchmarks in Appendix 8.1.
@@ 5.4 Efficiency
Figure 9: End-to-end runtime and memory consumption of Llava-Video-7B and 72B models with FrameFusion, using one and four NVIDIA A100-80GB GPUs for 7B and 72B models, respectively.
We evaluate the wall-clock speedup and GPU memory reduction of FrameFusion in Figure 9, across 7 to 72B model sizes and 32 to 256 frames. Results are averaged over 128 videos from the VideoMME benchmark. Additional results, including runtime and memory breakdowns, evaluations on more models and token budgets, and token reduction details, are provided in Appendix 8.2. Asymptotic complexity analysis are in Appendix 9.
Time. FrameFusion achieves end-to-end speedups of 1.6-3.6$\times$ with a 30% token budget. Compression stages in FrameFusion account for only 0.6-4% of the total runtime. The speedup grows as the number of frames increases, owing to the reduced $O((CN)^{2})$ complexity in attention computations. Since FrameFusion speedups the LLM computation in LVLMs, it yields greater speedups with the larger 72B model, where overhead such as video sampling and ViT encoding becomes proportionally less significant.
Memory. FrameFusion reduces GPU memory for KV-Cache and activations by the factor of token budget. As the number of frames grows and the dominance of model parameters decreases, memory savings continue to improve.
@@ 5.5 Ablation Study
We validate the impact of each observation and design choice in FrameFusion, showing their individual effectiveness with average scores on VideoNIAH, VideoMME (without subtitles), and NExT-QA benchmarks. We further analyze the effects of similarity computation strategies, distance metrics, similarity thresholds, and positional embeddings, as detailed in Appendix 8.3.
Design Choice 4.2. Instead of computing $N\times N$ cosine similarities with significant overhead, FrameFusion calculates only $N$ token similarities between corresponding visual tokens in adjacent frames. We compare our approach with two common alternative strategies: 1. Adj. token, which computes $N$ similarities between adjacent image patches (i.e., adjacent visual tokens). 2. Random, which calculates similarities for $N$ randomly selected token pairs. We compare the three similarity calculation strategies on three benchmarks at a 30% token budget. As shown in Table 2, our strategy achieves 11% and 8% higher average accuracies than the other two strategies. We also examine the choice of similarity calculation metrics by comparing cosine similarity with alternative distance metrics, including the inner product, Minkowski-2, and Minkowski-1 distance. Experiments show that cosine similarity yields 2-5% average score gains over the alternative metrics. The detailed numbers are presented in Appendix 8.3.2.
Design 4.2 VideoNIAH VideoMME NExT-QA Avg.
Random 60.0 59.0 56.6 58.5
Adj. token 64.0 59.9 56.6 60.2
Ours 76.5 61.3 56.8 64.9
@@
Table 2: Performance of different similarity calculation strategies with the same relative token budget of 30% on VideoNIAH, VideoMME, and NExT-QA.
Layer Rate VideoNIAH VideoMME NExT-QA Avg
Original 76.4 63.2 57.7 65.8
0 50.0% 76.2 62.7 57.4 65.4
1 52.0% 76.8 62.6 57.4 65.6
2 53.8% 76.4 62.0 57.3 65.2
12 87.5% 74.4 60.3 56.4 63.7
13 93.3% 64.2 57.9 55.6 59.2
14 100.0% 48.9 52.6 50.9 50.8
Table 3: Performance of Llava-Video-7B with cascaded token merging at different layers, where merging rates are adjusted to maintain an average token count (token budget) of 50% across all layers.
Design 4.2 VideoNIAH VideoMME NExT-QA Avg
Prune $\rightarrow$ merge 73.1 59.9 55.9 63.0
Merge $\rightarrow$ prune 73.3 60.9 56.6 63.6
@@
Table 4: Performance of the Llava-Video-7B model with different orders of merging and pruning.
@@
Design Choice 4.2. FrameFusion first merges at the initial layers, then prunes at subsequent layers. We evaluate the influence of the merging layer position. As shown in Table 3, delaying the merging to later layers requires a higher merging rate to meet the given relative token budget of 50%. The reduced high similarity values at deeper layers also causes significant performance drops compared with merging at shallower layers. Additionally, we examine the effect of the order of merging and pruning. As shown in Table 4, with a fixed token budget of 30% and the same number of tokens reduced in layers 1 and 2, merging before pruning achieves better performance compared to pruning before merging.
Design 4.2 VideoNIAH VideoMME NExT-QA Avg
Non-cascaded 74.9 60.0 55.4 63.4
Cascaded 76.0 62.8 57.5 65.4
Table 5: Performance of the Llava-Video-7B model with different orders of merging and pruning.
@@
Design Choice 4.2. FrameFusion adopts a cascaded merging strategy, where merged tokens remain reduced across layers to maximize efficiency. To evaluate the accuracy-efficiency trade-offs of cascaded merging, we also implement non-cascaded merging for comparison. Following previous works [26, 20, 43], this baseline performs the exact same merging strategy, but only on Key Value metrics in the attention module to avoid prominently removing any tokens. As a result, it leaves FFN computations unchanged. Its merging rate across all layers are set to be 30%. Our cascaded counterpart uses the same computation FLOPs, translating to an 84% relative token budget. As shown in Table 5, under the same FLOPs, FrameFusion outperforms the non-cascaded counterparts by 3% in average score. showing comparable performance to the original model.